# Relative monodromy of abelian logarithms for finite covers of universal families

### v2

@@ 7. $(vii)$
Thanks to the injective homomorphism obtained by Equation 29, we get $[\tau_{1,C}-\tau_{2,C}]_{C}=0$. By part $(vi)$ of 3.4, we also obtain $[\tau_{1}-\tau_{2}]_{S}=0$. By applying $\zeta_{3}$ we get $[\tau_{1}-\tau_{2}]_{C^{\prime}}=0$, which means that $\eta^{\prime}$ is injective.
Thus, by what we have just proven and by 3.8 the section $\sigma_{C}$ determines a cohomology class $[\sigma_{C}]_{C^{\prime}}$ in $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})\cong H^{1}(C^{\prime},\Lambda_{C^{\prime}})$. Again thanks to 3.8, we have that there exists a (possibly transcendental) section $\sigma^{\prime}:C^{\prime}\rightarrow\mathcal{A}^{\prime}|_{C^{\prime}}$ whose cohomology class $[\sigma^{\prime}]$ in $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})$ coincides with the class $[\sigma_{C}]_{C^{\prime}}$ defined above, i.e. $[\sigma^{\prime}]=[\sigma_{C}]_{C^{\prime}}$. Therefore, by 2.11 we obtain a section of the form
$\sigma_{C}-q_{C}^{*}(\sigma^{\prime})\in\Sigma^{\operatorname{an}}(C)$
that admits a logarithm. Notice that $\sigma_{C}-q_{C}^{*}(\sigma^{\prime})$ cannot be the zero-section, more precisely we can prove that $\sigma^{\prime}$ is a holomorphic section which is not algebraic. In fact, assume by contradiction that $\sigma^{\prime}$ is an algebraic section such that $q_{C}^{*}(\sigma^{\prime})=\sigma_{C}$. By part $(ii)$ of 3.4 and Equation 20 we have $\operatorname{Tr}(\sigma)|_{C^{\prime}}=\operatorname{Tr}(\sigma_{C})=\deg{q}\cdot\sigma^{\prime}$. By 2.1 $\sigma^{\prime}$ would be torsion, which is a contradiction.
Let us write
$\Sigma^{\operatorname{an}}(C)\ni\sigma_{C}-q_{C}^{*}(\sigma^{\prime})=\exp(f),$ (30)
@@
Now, by using properties of the operators $q_{C}^{*}$ and $\operatorname{Tr}$, we will see that $\sigma_{C}$ also has to admit a logarithm. Taking the trace of $f$, we can write
$f=\left(f-\frac{1}{\deg{q}}q_{C}^{*}\operatorname{Tr}(f)\right)+\frac{1}{\deg{q}}q_{C}^{*}\operatorname{Tr}(f)=l-l^{\prime},$
where
$l:=f-\frac{1}{\deg{q}}q_{C}^{*}\operatorname{Tr}(f),\quad\textrm{and}\quad l^{\prime}:=-\frac{1}{\deg{q}}q_{C}^{*}\operatorname{Tr}(f).$
[…]
$\sigma_{C}-q_{C}^{*}(\sigma^{\prime})=\exp(l)-\exp(l^{\prime}).$ (31)
Recalling that $\operatorname{Tr}(\sigma_{C})=0$ by part $(ii)$ of 3.4 and 2.1, taking traces on both sides of the previous equation and rearranging we obtain
00 0 $=\operatorname{Tr}(q_{C}^{*}(\sigma^{\prime})-\exp(l^{\prime}))=\deg{q}\cdot\sigma^{\prime}-\operatorname{Tr}(\exp(l^{\prime}))=$
$=\deg{q}\cdot\sigma^{\prime}-\exp(\operatorname{Tr}(l^{\prime})).$
[…]
@@ 3.3 Rank of relative monodromy
In this section we prove 1.2. We keep the same notations as above, in particular recall that $\rho:\pi_{1}(S(\mathbb{C}),s)\to\textrm{Mon}(\mathcal{A})$ denotes the monodromy representation associated to the abelian scheme $\phi:\mathcal{A}\to S$.
Theorem 3.11.
Let $\phi:\mathcal{A}\to S$ be an abelian scheme such that (up to a finite base change) the modular map $p:S\to\mathbb{A}_{g}$ is a finite surjective morphism. If $\sigma:S\to\mathcal{A}$ is a non-torsion section, then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{2g}$.
Proof.
Let us proceed by contradiction by assuming that $M_{\sigma}^{\textrm{rel}}$ is a submodule of $\mathbb{Z}^{2g}$ of rank $g^{\prime}$ with $g^{\prime}<2g$; notice that by 1.1 we have $g^{\prime}>0$. In other words, this means that for every $h\in H:=\ker{\rho}$ the logarithm $\log_{\sigma}$ is transformed by $h$ as
@@
where $v_{1},\ldots,v_{2g}\in\mathbb{Z}$ and $\omega_{1},\ldots,\omega_{2g}$ denote period functions. By 2.6 there exists $k\in\pi_{1}(S(\mathbb{C}),s)$ such that $\rho(k)\cdot M_{\sigma}^{\textrm{rel}}\nsubseteq M_{\sigma}^{\textrm{rel}}$, i.e. the action of $\rho(k)$ does not preserve the relative monodromy group. Therefore, we can fix $k\in\pi_{1}(S(\mathbb{C}),s)$ and $h\in H$ such that $\rho(k^{-1})\cdot(u_{1}^{h}\mu_{1}+\cdots+u_{g^{\prime}}^{h}\mu_{g^{\prime}})\notin M_{\sigma}^{\textrm{rel}}$; define $h^{\prime}:=k^{-1}hk$ and observe that $h^{\prime}\in H$ since $H\unlhd G$. If we look at the monodromy action of $h^{\prime}$ on the abelian logarithm we obtain
@@
Since $h^{\prime}\in H$, then we have $\rho(k^{-1})\cdot(u_{1}^{h}\mu_{1}+\cdots+u_{g^{\prime}}^{h}\mu_{g^{\prime}})\in M_{\sigma}^{\textrm{rel}}$, which contradicts the choice of $k$ and $h$. This concludes the proof.
@@ 4.1 Manin’s kernel theorem
A first application of the non-triviality of relative monodromy 1.1 is a strong version of Manin’s kernel theorem for an abelian scheme $\mathcal{A}\to S$ admitting a finite surjective map on some universal family without locally constant parts. The same result in the case of elliptic schemes was obtained in [9]. We refer to the Betti coordinates defined in Section 2.2 (see Equation 4).
@@
Let $\mathcal{A}\to S$ be an abelian scheme with a section $\sigma\in\Sigma(S)$. Assume that it admits a finite base change with a finite map on some universal family of abelian varieties without locally constant parts. The Betti coordinates $\beta_{\sigma,i}$ with $i=1,\ldots,2g$ are globally defined functions on $S(\mathbb{C})$ if and only if $\sigma$ is a torsion section. In that case, the Betti coordinates are rational constants.
@@
Clearly, if $\sigma$ is a torsion section the claim follows by definition. Conversely, if the Betti coordinates are globally defined on $S(\mathbb{C})$ then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is trivial. By 1.1 the section must be torsion.
@@ 4.2 Algebraic independence of periods and logarithms
Our theorem about the rank of the relative monodromy group 1.2 gives a different proof of a result due to André [1, Theorem 3] about the algebraic independence of logarithms under our hypothesis on the abelian scheme $\mathcal{A}\to S$. Let’s briefly describe the setting.
Recall that the monodromy action of the fundamental group $\pi_{1}(S(\mathbb{C}),s)$ on periods and the logarithm of a section $\sigma$ induces a representation $\theta_{\sigma}:\pi_{1}(S(\mathbb{C}),s)\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z})$ described in Equation 12. Consider the projection $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))\to\textrm{Sp}_{2g}(\mathbb{Z})$, whose image is the monodromy group $\textrm{Mon}(\mathcal{A})$. By general theory, the Zariski-closure $\overline{\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))}$ in $\textrm{SL}_{2g+1}$ is the differential Galois group of the Picard-Vessiot extension of $\mathbb{C}(S)$ generated by the coordinates of $\omega_{1},\ldots,\omega_{2g},\log_{\sigma}$ and their directional derivatives along a tangent vector field $\partial$ with respect to the Gauss-Manin connection. Clearly the homomorphism $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))\to\textrm{Sp}_{2g}(\mathbb{Z})$ extends to an algebraic group homomorphism
@@
Here the aforementioned result [1, Theorem 3] in the case where the abelian scheme is endowed (up to a base change) with a finite surjective morphism on a suitable universal family.
@@
Let $\mathcal{A}\to S$ be an abelian scheme admitting (up to a base change) a finite surjective morphism $S\to\mathbb{A}_{g}$, where $\mathfrak{A}_{g}\to\mathbb{A}_{g}$ is a universal family of abelian varieties without locally constant parts. If $\sigma\in\Sigma(S)$ is a non-torsion section, the kernel $\ker{\xi}$ is isomorphic to $\mathbb{G}_{a}^{2g}$. In particular, the coordinates of $\log_{\sigma}$ have transcendence degree $g$ over $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g})$, where $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g})$ denotes the extension of $\mathbb{C}(S)$ generated by the coordinates of periods.
@@
By 1.2 the kernel of the morphism $\xi|_{\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))}$ is isomorphic to $\mathbb{G}_{a}^{2g}$, which implies the first part of the statement. In particular, the transcendence degree of the extension generated by the coordinates of $\log_{\sigma}$ and their directional derivatives over $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g},\partial\omega_{1},\ldots,\partial\omega_{2g})$ is $2g$, where we are fixing a tangent vector field $\partial$. The conclusion follows.

### v3

@@ 7. $(vii)$
Thanks to the injective homomorphism obtained by Equation 30, we get $[\tau_{1,C}-\tau_{2,C}]_{C}=0$. By part $(vi)$ of 3.6, we also obtain $[\tau_{1}-\tau_{2}]_{S}=0$. By applying $\zeta_{3}$ we get $[\tau_{1}-\tau_{2}]_{C^{\prime}}=0$, which means that $\eta^{\prime}$ is injective.
Thus, by what we have just proven and by 3.9 the section $\sigma_{C}$ determines a cohomology class $[\sigma_{C}]_{C^{\prime}}$ in $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})\cong H^{1}(C^{\prime},\Lambda_{C^{\prime}})$. Again thanks to 3.9, we have that there exists a (possibly transcendental) section $\sigma^{\prime}:C^{\prime}\rightarrow\mathcal{A}^{\prime}|_{C^{\prime}}$ whose cohomology class $[\sigma^{\prime}]$ in $H^{1}(\pi_{1}(C^{\prime},s^{\prime}),\mathbb{Z}^{2g})$ coincides with the class $[\sigma_{C}]_{C^{\prime}}$ defined above, i.e. $[\sigma^{\prime}]=[\sigma_{C}]_{C^{\prime}}$. Therefore, by 2.9 we obtain a section of the form $\sigma_{C}-q_{C}^{*}(\sigma^{\prime})\in\Sigma^{\operatorname{an}}(C)$ that admits a logarithm; let us write
$\sigma_{C}-q_{C}^{*}(\sigma^{\prime})=\exp(f),$ (31)
@@
Now, by using properties of the operators $q_{C}^{*}$ and $\operatorname{Tr}$, we will get a contradiction. Applying first the trace operator and then the pullback operator we obtain:
$q^{*}_{C}\operatorname{Tr}(\sigma_{C})-\deg{q}\cdot q^{*}_{C}(\sigma^{\prime})=\exp(q^{*}_{C}\operatorname{Tr}(f)).$
By using Equation 31, we finally obtain
$q^{*}_{C}\operatorname{Tr}(\sigma_{C})-\deg{q}\cdot\sigma_{C}=\exp(q^{*}_{C}\operatorname{Tr}(f)-\deg{q}\cdot f).$
Thus, the section $q^{*}_{C}\operatorname{Tr}(\sigma_{C})-\deg{q}\cdot\sigma_{C}$ is algebraic and admits a global logarithm. Notice that
$q^{*}_{C}\operatorname{Tr}(\sigma_{C})-\deg{q}\cdot\sigma_{C}=(q^{*}\operatorname{Tr}(\sigma)-\deg{q}\cdot\sigma)_{|_{C}}.$
By part $(vi)$ of 3.6, the section $q^{*}\operatorname{Tr}(\sigma)-\deg{q}\cdot\sigma$ is also an algebraic section admitting a global logarithm on $S(\mathbb{C})$. By 3.3 and [22, Proposition 2.10] the base $S$ can be thought to be maximal. By 2.8 we must have $\sigma=q^{*}\operatorname{Tr}(\sigma)/\deg{q}$, which leads to a contradiction with the fact that $\sigma$ is a ramified section.
Remark 3.11.
Note that, under the hypothesis of generic finiteness of the modular map, from 2.6 and 1.2 it follows immediately that a torsion section $\sigma\colon S\to\mathcal{A}$ is unramified. When $S$ is a curve this follows also from [31, Theorem 1].
@@ 3.3 Rank of relative monodromy
Here, we prove 1.3.
We keep the same notations as above, in particular recall that $\rho:\pi_{1}(S(\mathbb{C}),s)\to\textrm{Mon}(\mathcal{A})$ denotes the monodromy representation associated to the abelian scheme $\phi:\mathcal{A}\to S$, which we are assuming to induce an irreducible action of $\textrm{Mon}(\mathcal{A})$ on the lattice of periods. In what follows we identify the $\mathbb{Z}$-module $M_{\sigma}^{\textrm{rel}}$ with the corresponding $\mathbb{Z}$-submodule of the lattice of periods.
Proof of 1.3
Let us proceed by contradiction by assuming that $M_{\sigma}^{\textrm{rel}}$ is a submodule of $\mathbb{Z}^{2g}$ of rank $g^{\prime}$ with $g^{\prime}<2g$. By the non-triviality of $M_{\sigma}^{\textrm{rel}}$ proved in 1.2 we have $g^{\prime}>0$. In other words, this means that for every $h\in H:=\ker{\rho}$ the logarithm $\log_{\sigma}$ is transformed by $h$ as
@@
where $v_{1},\ldots,v_{2g}\in\mathbb{Z}$ and $\omega_{1},\ldots,\omega_{2g}$ denote period functions. Since we are assuming that the action of $\textrm{Mon}(\mathcal{A})$ is irreducible, there exists $k\in\pi_{1}(S(\mathbb{C}),s)$ such that $\rho(k)\cdot M_{\sigma}^{\textrm{rel}}\nsubseteq M_{\sigma}^{\textrm{rel}}$, i.e. the action of $\rho(k)$ does not preserve the relative monodromy group. Therefore, we can fix $k\in\pi_{1}(S(\mathbb{C}),s)$ and $h\in H$ such that $\rho(k^{-1})\cdot(u_{1}^{h}\mu_{1}+\cdots+u_{g^{\prime}}^{h}\mu_{g^{\prime}})\notin M_{\sigma}^{\textrm{rel}}$; define $h^{\prime}:=k^{-1}hk$ and observe that $h^{\prime}\in H$ since $H\unlhd G$. If we look at the monodromy action of $h^{\prime}$ on the abelian logarithm we obtain
@@
Since $h^{\prime}\in H$, then we have $\rho(k^{-1})\cdot(u_{1}^{h}\mu_{1}+\cdots+u_{g^{\prime}}^{h}\mu_{g^{\prime}})\in M_{\sigma}^{\textrm{rel}}$, which contradicts the choice of $k$ and $h$. This concludes the proof. ∎
Remark 3.12.
When $\mathcal{A}\to S$ is simple and $S$ is a curve, maximal variation in moduli is equivalent to not having fixed part. Hence our 1.2 holds for all ramified sections of an abelian scheme with no fixed part over a curve $S$.
An interesting case where our result holds, is given by the family of hyperelliptic curves described in [3, Section 7].
The hypothesis on the irreducibily of the action of $\text{Mon}(\mathcal{A})$ appearing in 1.3 is satisfied when $p(S)$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$.
3.4 Relative monodromy of non-torsion sections
So far all the results are proved for ramified sections, which are a subset of all non-torsion sections (see 3.11). We can extend our main results to all non-torsion sections under the assumption that the modular map $p$ is generically finite onto a subvariety $\mathcal{G}\subseteq\mathbb{A}_{g}$, where the unramified base changes of $\mathcal{G}$ have finite Mordel-Weil group.
Proposition 3.13.
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to\mathcal{G}\subseteq\mathbb{A}_{g}$ is generically finite, where the unramified base changes of $\mathcal{G}$ have finite Mordel-Weil group. If $\sigma:S\to\mathcal{A}$ is a non-torsion section, then $M_{\sigma}^{\textrm{rel}}$ is non-trivial. If, in addition, the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ is irreducible, then $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{2g}$.
Proof.
Since the unramified base changes of $\mathcal{G}$ have finite Mordell-Weil group, each unramified section is torsion. The result immediately follows from 1.2 and 1.3.
The condition about the Mordell-Weil group holds true when $\mathcal{G}$ is a Kuga family (see [26, Main Theorem]). In particular, $\mathcal{G}$ may be chosen to be a totally geodesic (i.e. weakly-special) subvariety of $\mathbb{A}_{g}$. Notice that, if we pick $\mathcal{G}=\mathbb{A}_{g}$ we obtain the complete version of 1.4: in fact we don’t need to distinguish the cases when the image of the section $\sigma$ is contained in proper group subschemes since having a generically finite map on a universal family prevents our abelian scheme from having any non-trivial proper subscheme; for the same reason, in this case the scheme automatically does not have fixed part.
3.5 The case of thin Monodromy of periods
In [2] André uses some Hodge-theoretical methods to address the problems presented in this paper. These are his hypotheses:
  • $(i)$
    $\mathcal{A}\to S$ has no fixed part, even after pull-back to a finite étale cover of $S$,
  • $(ii)$
    the image of the section is not contained in any proper group-subgroup,
  • $(iii)$
    the monodromy group $\textrm{Mon}(\mathcal{A})$ is not thin, i.e. it is of finite index in the integral points of its Zariski closure,
He shows that $M_{\sigma}^{\textrm{rel}}\cong\mathbb{Z}^{2g}$. It would be interesting to compare his results with ours.
Regarding the hypothesis $(iii)$, André observes that it is not always satisfied and in general we don’t know how big is $M_{\sigma}^{\textrm{rel}}$ in such case. To this regard, we provide below an explicit construction showing a case where $M_{\sigma}^{\textrm{rel}}$ is of maximal rank even though $\textrm{Mon}(\mathcal{A})$ is thin.
Remark 3.14.
Notice that a priori, our hypotheses in 1.3 don’t exclude the case when $\textrm{Mon}(\mathcal{A})$ is thin.
Example 3.15.
Let us denote by $\mathcal{L}\to S$ the Legendre elliptic scheme, where $S:=\mathbb{P}_{1}-\{0,1,\infty\}$. Consider the product
$(\phi_{1},\phi_{2}):\mathcal{L}\times\mathcal{L}\to S\times S,$
where we use the coordinates $(x,y)$ in $S\times S$. We let $X$ be the line in $S^{2}$ defined by $x+y=2$, so this line is isomorphic under the first projection to $B:=\mathbb{P}_{1}-\{0,1,2,\infty\}$. By projection from $X$ to $B$, we may define a scheme $\phi:\mathcal{A}\to B$ over $B$ whose fibers are of the type $\phi^{-1}(\lambda)=\mathcal{L}_{\lambda}\times\mathcal{L}_{2-\lambda}$.
The abelian scheme $\phi:\mathcal{A}\to B$ may be interpreted as the fiber product of two elliptic schemes $\mathcal{E}_{1}\times_{B}\mathcal{E}_{2}\to B$, where $\mathcal{E}_{1}\to B$ is the Legendre scheme associating $\mathcal{L}_{\lambda}\to\lambda$ and $\mathcal{E}_{2}\to B$ is the Legendre scheme associating $\mathcal{L}_{\lambda}\to 2-\lambda$. Note that the two elliptic schemes are not isogenous, since they have a different set of bad reduction.
The fundamental group $G$ of $B(\mathbb{C})$ is free on three generators, and the monodromy representation on the periods of $\phi$ gives a homomorphism $\rho:G\to\Gamma_{2}\times\Gamma_{2}$, where $\Gamma_{2}$ is the subgroup of $\textrm{SL}_{2}(\mathbb{Z})$ generated by
$\begin{pmatrix}1&2\\ 0&1\end{pmatrix},\qquad\begin{pmatrix}1&0\\ 2&1\end{pmatrix}.$
We denote the monodromy group of periods as $\textrm{Mon}(\mathcal{A}):=\rho(G)$. Since $\phi:\mathcal{A}\to B$ is the product of two non-isogenous elliptic schemes, then the Zariski closure of $\textrm{Mon}(\mathcal{A})$ is the whole $\textrm{SL}_{2}\times\textrm{SL}_{2}$ [9, Theorem 5.3.10].
Notice that $\textrm{Mon}(\mathcal{A})$ is generated by three elements, and the same holds for its projection on the abelianized group of $\Gamma_{2}\times\Gamma_{2}$, which is $\mathbb{Z}^{4}$. In particular, the projection in the abelianized group must have infinite index. Therefore, the group $\textrm{Mon}(\mathcal{A})$ is not of finite index in the integral points of its Zariski-closure: in other words, the group $\textrm{Mon}(\mathcal{A})$ is thin.
The abelian scheme $\phi\colon\mathcal{A}\to B$ has trivial Mordell-Weil rank. Let us consider the map $\sigma:B\to\mathcal{A}$ defined by the matrices
$\sigma(\lambda):=\left(\left(2,\sqrt{2(2-\lambda)}\right),\left(3,\sqrt{6(1+\lambda)}\right)\right).$
This map is a multisection of the abelian scheme $\phi$, but it becomes a section whether we consider the base change induced by the degree $4$ field extension $\mathbb{C}(B)\subseteq\mathbb{C}(B)(\psi,\mu)$ where $\psi^{2}+\mu^{2}=3$ and $\mu^{2}=1+\lambda$. Denote by $\phi^{\prime}:\mathcal{A}^{\prime}\to B^{\prime}$ the base change we just described. The monodromy of periods $\textrm{Mon}(\mathcal{A}^{\prime})$ is a finite index subgroup of $\textrm{Mon}(\mathcal{A})$: in particular, it is Zariski dense in $\textrm{SL}_{2}\times\textrm{SL}_{2}$ and it has infinite index in $\Gamma_{2}\times\Gamma_{2}$; so $\textrm{Mon}(\mathcal{A}^{\prime})$ is still thin. The morphism $\sigma$ is a section of $\phi^{\prime}$ whose image is not contained in any proper group subscheme, thus the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{4}$ by [33, Theorem 4.11].
@@ 4.1 Manin’s kernel theorem
A first application of the non-triviality of relative monodromy is a strong version of Manin’s kernel theorem under the hypotheses of our 3.13. In the case of elliptic schemes the same result was obtained in [10]. We refer to the Betti coordinates defined in Section 2.2 (see Equation 5).
@@
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to\mathcal{G}\subseteq\mathbb{A}_{g}$ is generically finite, where the unramified base changes of $\mathcal{G}$ have finite Mordel-Weil group. Let $\sigma:S\to\mathcal{A}$ be a section. If the Betti coordinates $\beta_{\sigma,i}$ with $i=1,\ldots,2g$ are globally defined functions on $S(\mathbb{C})$, then $\sigma$ is a torsion section. In that case, the Betti coordinates are rational constants.
@@
If the Betti coordinates are globally defined on $S(\mathbb{C})$ then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is trivial. The claim then follows by the non-triviality part of 3.13. ∎
@@ 4.2 Algebraic independence of periods and logarithms
Our theorem about the rank of the relative monodromy group 1.3 gives a different proof of a result due to André [1, Theorem 3] about the algebraic independence of logarithms under our hypothesis on the abelian scheme $\mathcal{A}\to S$. Let’s briefly describe the setting.
Recall that the monodromy action of the fundamental group $\pi_{1}(S(\mathbb{C}),s)$ on periods and the logarithm of a section $\sigma$ induces a representation $\theta_{\sigma}:\pi_{1}(S(\mathbb{C}),s)\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z})$ described in Equation 13. Consider the projection $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))\to\textrm{Sp}_{2g}(\mathbb{Z})$, whose image is the monodromy group $\textrm{Mon}(\mathcal{A})$. By general theory, the Zariski closure $\overline{\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))}$ in $\textrm{SL}_{2g+1}$ is the differential Galois group of the Picard-Vessiot extension of $\mathbb{C}(S)$ generated by the coordinates of $\omega_{1},\ldots,\omega_{2g},\log_{\sigma}$ and their directional derivatives along a tangent vector field $\partial$ with respect to the Gauss-Manin connection. Clearly the homomorphism $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))\to\textrm{Sp}_{2g}(\mathbb{Z})$ extends to an algebraic group homomorphism
@@
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to p(S)\subseteq\mathbb{A}_{g}$ is generically finite. If $\sigma:S\to\mathcal{A}$ is a ramified section and the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ is irreducible, then the kernel $\ker{\xi}$ is isomorphic to $\mathbb{G}_{a}^{2g}$. In particular, the coordinates of $\log_{\sigma}$ have transcendence degree $g$ over $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g})$, where $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g})$ denotes the extension of $\mathbb{C}(S)$ generated by the coordinates of periods.
@@
By 1.3 the kernel of the morphism $\xi|_{\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))}$ is isomorphic to $\mathbb{G}_{a}^{2g}(\mathbb{Z})$, which implies the first part of the statement. In particular, the transcendence degree of the extension generated by the coordinates of $\log_{\sigma}$ and their directional derivatives over $\mathbb{C}(S)(\omega_{1},\ldots,\omega_{2g},\partial\omega_{1},\ldots,\partial\omega_{2g})$ is $2g$, where we are fixing a tangent vector field $\partial$. The conclusion follows.